3: Determine the reactions at A and C. (20 pts) w=5 lb/ft W2=10 lb/ft + W2=5 lb/ft 2. 3 А С 4 - 7 - 5
Answer #1

10 16 olb 1.333 2.666 2 He 7 RA 4. EFx = 0 -He 5X10= IHc -10 lb 09 E Fyo (584) – (5x5x4) + Ve = 0 RA RAT Nc 20+10 = 30 EMA=0